OpenStudy (anonymous):

desperate help needed :) find the mass of this solute - 200ml of 1.5M sodium hydroxide- explanation plz :)

OpenStudy (frostbite):

Right, from me you always get an explanation: We need 2 very simple equations here: \[\LARGE C=\frac{ n }{ V }\] \[\LARGE n=\frac{ m }{ M }\] Here is: C: the concentration [M = mol/L] n: the amount of substance [mol] V: the volume [L] M: the molecular mass [g/mol] m: the mass [g]

OpenStudy (anonymous):

12 g

OpenStudy (frostbite):

We notice we have a concentration and a volume. from that we can calculate the amount of substance: \[\LARGE n=CV\] When we have that number we can simply use it in the other equation and solve for the mass: \[\LARGE CV=\frac{ m }{ M } \to m=CVM\]

OpenStudy (anonymous):

thankyou :)

OpenStudy (frostbite):

Oh yea and M(NaOH) is almost equal to 40 gram / mol... but you know how to find it your self?

OpenStudy (anonymous):

you mean to find the mol about of sodium hydroxide yeah :) yes

OpenStudy (frostbite):

No I mean the molecular mass: \[\large M(NaOH)=M(Na)+M(O)+M(H )\approx 40 \frac{ g }{ mol }\]

OpenStudy (anonymous):

yes i meant that all good thanks

OpenStudy (frostbite):

Perfect, not problem at all.

OpenStudy (anonymous):

:)